OpenStudy (anonymous):

TOPIC: Indices and logarithms. Simplify 3^n+2 - 5(3^n) - 6 (3^n-1)

OpenStudy (anonymous):

Hi :3

OpenStudy (anonymous):

5*3^n

OpenStudy (callisto):

I assume your question is like this 3^(n+2) - 5(3^n) - 6 [3^(n-1)] Then, 3^(n+2) - 5(3^n) - 6 [3^(n-1)] = (3^2)3^(n) - 5(3^n) - 6(3^-1) 3^(n) = 9(3^n) - 5(3^n) - 2(3^n) =2(3^n)

OpenStudy (anonymous):

thanks, you're a big helper.

OpenStudy (callisto):

welcome :) but i'm not a 'big' helper, I'm just a mini one :)

OpenStudy (anonymous):

ahah, okay